OpenStudy (anonymous):

Could someone please explain this? Solve by completing the square: x² + 10x = 18 a)–11.56, 1.56 b)11.56, 1.56 c)–11.56, –1.56 d)11.56, –1.56

OpenStudy (amorfide):

completing the square: you will do (x+b/2)²-(b/2)²=18 add -(b/2)² to both sides now square root both sides to get x+b/2 on the left hand side this will give x+b/2= plus or minus root of whatever is on the right -b/2 on both sides x=__________

OpenStudy (anonymous):

I did that...I didn't get any of the answers listed!